OpenStudy (anonymous):

what are the vertex and axis of symmetry of the equation y=-2x^2+8x-18

OpenStudy (anonymous):

https://www.mathway.com/

OpenStudy (anonymous):

i tried it dont help

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

would you help me i dont know how to find it

OpenStudy (anonymous):

give me a minute

OpenStudy (jack1):

vertex is the high point, yeah?

OpenStudy (anonymous):

look it up on google n the answer is there it was already asked on openstudy hope that helped.

OpenStudy (anonymous):

i am with you jack

OpenStudy (anonymous):

y=−2x2+8x−18 y+18−8=−2(x2−4x+4) y+10=−2(x−2)2

OpenStudy (jack1):

cool, so ur looking for ur turning point ur equation's: y=-2x^2+8x-18

OpenStudy (anonymous):

yes

OpenStudy (jack1):

y = -2x^2+8x-18 so ur derivative's (y') y' =-4x + 8

OpenStudy (jack1):

you've done derivatives yet, yeah?

OpenStudy (anonymous):

no not yet

OpenStudy (anonymous):

i am with you though

OpenStudy (anonymous):

i have answer choices would that help you in away

OpenStudy (jack1):

dam...ok the short of derivatives is: bring the power down, multiply by co-efficient, then put back up and minus 1 so if u had: y = 10x^4 -12x^3 +9x^2 + 2x + 19 derivative would be y' = 10*4 x^4-1 - 12*3 x^3-1 + 9*2 x*2-1 + 2*1 x^1-1 + 0 so y' = 40x^3 - 36x^2 + 18x*1 + 2x^0 y' = 40x^3 - 36x^2 + 18x*1 + 2 make sense?

OpenStudy (jack1):

either way: you're derivative is the slope of the curve at any x or y point along it so y' =-4x + 8 when slope = 0, its a turning point so y' =-4x + 8 0 =-4x + 8 -8 =-4x x = 2 your vertex occurs at x = 2 whats the y value at this point? @jamesx3rd ?

OpenStudy (anonymous):

yeah you times it the minus the power down by one

OpenStudy (anonymous):

i see how you get the answers i am with i am actualy learning this from you

OpenStudy (anonymous):

what do you do with the x=2 now though

OpenStudy (anonymous):

were did the y go?

OpenStudy (jack1):

hall good dude, happy ur learning y = -2x^2+8x-18 if x = 2 y = -2x^2+8x-18 y = -2(2^2) +8(2) -18 y = -2*4 + 8*2 - 18 y = -8 + 16 - 18 y = -10 so when x = 2, y = -10 so vertex occurs at (2, -10)

OpenStudy (anonymous):

my answer choices are: Vertex: (2,-10) axis of symmetry x=2 i think this one is right Vertex: (2,-10) axis of symmetry x =-10 Vertex: (-2,-10) axis of symmetry x=-2 Vertex (-2,10) axis of symmetry x=-2

OpenStudy (anonymous):

you plug the 2 in the problem and got that -10 right

OpenStudy (jack1):

may be helpful in the meantime, until u get onto derivatives in a big way : The standard equation of a parabola is y = ax2 + bx + c. But the equation for a parabola can also be written in "vertex form": y = a(x – h)2 + k In this equation, the vertex of the parabola is the point (h, k). The coefficient of x here is –2ah. This means that in the standard form, y = ax2 + bx + c, the expression -b/2a gives the x-coordinate of the vertex.

OpenStudy (anonymous):

so was the first one right

OpenStudy (jack1):

and yep, a is correct

OpenStudy (anonymous):

would you help me on one more problem ill make another post

OpenStudy (jack1):

yep ill follow

OpenStudy (anonymous):

thanks